# The Emerged Security and Privacy of LLM Agent: A Survey with Case Studies

### v1

@@ 5.3.1. Defense on Knowledge Poisoning
(Baracaldo et al., 2017) proposes a new method for detecting and filtering poisonous data in the training sets of supervised learning models. It specifically utilizes data provenance to identify groups of data with a high correlation in their likelihood of being poisoned. This innovative approach aids in the effective identification and removal of malicious data.  (Yan et al., 2023) presents ParaFuzz, a novel framework for detecting poisoned samples at test time in large language models (LLMs), leveraging the interpretability of model predictions. The effectiveness of PARAFUZZ heavily depends on the specific prompts used with ChatGPT, which is employed to ensure high-quality paraphrasing. To optimize the detection process, the study adopts fuzzing to develop precise paraphrase prompts. These prompts are designed to effectively neutralize backdoor triggers while preserving the semantic integrity of the text.
There is still a significant gap in research focused on developing efficient defense strategies to protect LLMs from knowledge poisoning attacks (Das et al., 2024). Furthermore, empirical evidence indicates that LLMs are increasingly susceptible to these attacks. Current defense mechanisms, such as filtering data or reducing model capacity, provide only limited protection and often result in decreased test accuracy (Wan et al., 2023b).
@@
5.3.2. Defense on Functional Manipulation
Given the emergence of Functional Manipulation as a new risk associated with the deployment of LLM agents, research on this specific threat remains limited. Thus, proactive security measures are essential. When using third-party LLM agents, it is crucial to protect personal privacy and be wary of excessive personal data requests by third parties. Users should limit data sharing, especially avoiding sensitive or personally identifiable information during interactions with LLM agents. Additionally, understanding and utilizing the data protection settings offered by LLM agents is vital. Adjusting privacy settings helps control what data can be collected and processed. Choosing providers with a strong reputation and transparency is also recommended, as these providers should have clear data usage and privacy protection policies along with a robust security track record (Zhang et al., 2024a).
Furthermore, to address the challenges posed by Functional Manipulation, the introduction of the ToolEmu (Ruan et al., 2024) framework represents a significant advancement. This framework employs a language model to emulate tool execution, which allows for extensive and scalable testing of LM agents across diverse scenarios and toolsets. Coupled with an LM-based automatic safety evaluator, ToolEmu facilitates the identification and quantification of risks by examining potential failures and subsequent consequences. This method provides a dynamic alternative to traditional static sandbox evaluations, enhancing the ability to detect and mitigate high-stakes, long-tail risks effectively.
Additionally,  (Anderljung et al., 2023) proposes an initial set of safety standards as an essential first step in industry self-regulation. These standards include pre-deployment risk assessments, external reviews of model behavior, the use of risk assessments to inform deployment decisions, and monitoring and responding to new information about model functionality post-deployment. This approach contributes valuable insights to the broader discussion on balancing public safety risks with the benefits of innovation in AI development.
@@ 5.3.3. Defense on Output Manipulation
Inspired by human recursive thinking in the Avalon game,  (Wang et al., 2023b) introduces Recursive Contemplation (ReCon), a framework designed to enhance LLMs’ ability to detect and counter deceptive information. ReCon employs formulation, which generates initial thoughts and speech, and refinement, which improves these outputs. It also includes two perspective transitions, aiding LLMs in understanding others’ mental states and how others perceive their own mental states.
Additionally,  (Xu et al., 2023a) has developed a benchmarking framework called MAgIC, designed to evaluate LLMs in multi-agent environments. It utilizes games and game theory scenarios to test models on reasoning, cooperation, and adaptability. The research employs Probabilistic Graphical Modeling (PGM) to enhance models’ capabilities in handling complex social interactions.
@@ 6. Future Trends and Discussion
With the continuous advancements in LLM agents, these agents have become capable of effectively interacting with users through complex observations, reasoning, and task execution, demonstrating broad application prospects across multiple domains. Particularly with the development of Multimodal Large Language Model (MLLM) agents, LLM agents can now process various data types, including text, images, and audio, significantly expanding their application scope. Moreover, by incorporating Large Language Model Multi-Agent (LLM-MA) systems, different LLM agents can collaborate to accomplish more complex tasks. The integration of these technologies will contribute to building more intelligent and efficient systems. However, the widespread application of these advanced technologies also introduces significant challenges related to privacy and security. Through discussions on future trends, our aim is to provide insights for researchers, developers, and policymakers on how to optimize these technologies and overcome related challenges.
@@ 6.1.1. The Development of MLLM Agent
Recent advancements in LLMs have significantly surpassed traditional boundaries of language processing. These models now incorporate supplementary components such as instruction, interface, tools, knowledge, and memory, evolving into intelligent LLM agents that demonstrate expanded reasoning and expertise. Research studies  (Yang et al., 2023a; Wu et al., 2023) indicate efforts to bridge the gap between language models and multimodal tools, with intelligent agents like Visual ChatGPT  (Wu et al., 2023) and MMREACT  (Yang et al., 2023a) employing sophisticated prompt engineering techniques to achieve this target. Such efforts have given rise to the field of Multimodal Large Language Models (MLLMs). The general architecture of the MLLM is depicted in Figure 12
@@
MLLMs are based on LLMs and enhanced with the capability to receive, reason, and output multimodal information. By integrating various data modalities, such as text, image, audio and video, these models are not only capable of understanding information from a single modality but can also process and interpret across modalities, thus achieving a comprehensive understanding of complex information (Yin et al., 2023a). The application of MLLMs has extended to several other fields, including medical image analysis (Zhang et al., 2023a; Moor et al., 2023) and document processing (Hu et al., 2024b; Liu et al., 2024c).
Moreover, the development of multimodal agents based on MLLMs, such as embodied agents (Huang et al., 2024c) and graphical user interface agents (Wang et al., 2024b), has further enhanced these models’ interactive capabilities in physical environments. These agents, utilizing MLLMs as planners and following natural language instructions to navigate and interact effectively in the real world, are not only designed to understand and generate information but are also equipped with essential skills such as perception, reasoning, planning, and execution. This enables them to operate effectively in complex real-world environments  (Xie et al., 2024).
With the advent of MLLM agents, the potential for achieving Artificial General Intelligence (AGI) has become more feasible, leading to significant advancements in Embodied AI. The ability of agent robots to understand and respond to human commands is crucial, especially in service-oriented tasks. The substantial progress in MLLMs has equipped them with the ability to comprehend and generate natural human instructions effectively. This progress could enable robots to learn user preferences and provide services that closely mimic human interaction.
@@ 6.1.2. The Security and Privacy Research on MLLM Agent
The development of embodied agents capable of interacting with the real world becomes a highly active area of research. However, MLLM agents also present several security vulnerabilities, one of which is the phenomenon of multimodal hallucinations.
@@
Unlike language hallucinations, multimodal hallucinations refer to the phenomenon where the output descriptions generated by MLLMs are inconsistent with the actual content of images  (Yin et al., 2023b), as shown in Figure 13. These phenomena manifest in two primary forms (Lee et al., 2024): one involves generated content that includes objects which are inconsistent with or absent from the target image  (Zhai et al., 2024; Liu et al., 2024a); the other, a more complex form, encompasses holistic misrepresentations of entire scenes or environments (Sun et al., 2023).
Current methods to reduce these hallucinations encompass several approaches, such as utilizing self-feedback with visual cues to enhance model accuracy  (Lee et al., 2024), employing instruction-tuning techniques to refine the model’s response to respond to human instructions  (Liu et al., 2024a). implementing error-correction processes that identify and rectify hallucinations within the generated text  (Yin et al., 2023b). Despite these efforts, significant challenges remain, requiring a sophisticated ability to distinguish between accurate and hallucinatory outputs, along with improvements in training approaches to boost the reliability of the outputs.
Similar to LLM agents, MLLM agents can be susceptible to crafted attacks (Qi et al., 2023; Bagdasaryan et al., 2023; Shayegani et al., 2023). These agents may be maliciously manipulated to produce biased or undesirable responses. However, research in this area is still in its early stages. Therefore, enhancing the safety of these MLLM agents is an essential focus of ongoing research. Improvements in MLLM agents safety will involve developing robust mechanisms to detect and mitigate these vulnerabilities, ensuring that MLLM agents can function reliably and securely in diverse applications. Such advancements are crucial for the broader adoption and ethical deployment of AI technologies in real-world environment.
@@ 6.2.1. The Development of LLM-MA System
LLM agents exhibit advanced reasoning and planning capabilities, approaching human-like levels of decision-making and interaction. These agents are adept at perceiving their environments, making informed decisions, and executing actions based on complex contexts (Yao et al., 2024).
@@
The application of LLM-MA systems spans across various fields, broadly categorized into two main types: problem solving and world simulation (Guo et al., 2024). For problem-solving applications, such as multi-robot systems (Mandi et al., 2023) and software development (Du et al., 2024), these systems enable interactions among diverse agents. This collaborative capability effectively solves complex real-world problems, mirroring the cooperative nature of human group work in tackling multifaceted challenges. On the other hand, world simulation encompasses applications such as society simulations (Park et al., 2023b) and game simulation (Xu et al., 2024). The case studies parts presented in this paper illustrates the application of world simulation to depict the threats faced by LLM agents and their impacts, presenting one of the many facets of LLM-MA systems utilization.
@@ 6.2.2. The Security and Privacy Research on LLM-MA System
As research on LLM-MA system increases rapidly, numerous challenges have emerged. Each agent in a multi-agent system may need to access and process sensitive data, and even execute code. This has sparked discussions on the security and privacy concerns related to multi-agent systems.
Each agent within a multi-agent system may need to access and process sensitive data, and even execute code. Moreover, due to the intercommunication and interconnection between agents, security issues originating from a single agent can have profound and amplified effects in a multi-agent scenario. This has intensified the need for focused discussions on security and privacy issues in multi-agent environments.
The issue of hallucination, where agents generate outputs based on incorrect or fabricated information, represents a significant challenge for both LLMs and LLM Agents. This problem becomes even more complex in a multi-agent context due to the interconnected nature of these agents and their frequent communication. Misinformation from one agent can be accepted and further propagated by others within the network, leading to a cascade of misinformation. To mitigate this issue, it is crucial to correct errors at the individual agent level and also to manage the flow of information between agents, thereby preventing the spread of inaccurate information throughout the entire system (Guo et al., 2024).
Furthermore, the capability of LLM multi-agent systems to interact with files and execute code offers extensive possibilities for their application. However, the presence of potentially malicious LLM agents within the system poses significant risks. In one case, these agents may operate in a passive listening mode, where they receive information shared by other agents to perform tasks, but at the same time, they leak confidential information to attackers deliberately. In the other case, malicious LLM agents may engage in a active communication mode, spreading virus-infected files, phishing messages, or other malicious codes, attempting to attack or disrupt other agents within the system. To mitigate this risk, incorporating human feedback and user authorization for each step can help reduce these threats. This necessitates designing the system with robust security measures to prevent unauthorized access or misuse. One effective approach is the implementation of a state-less oracle agent, which can monitor each sensitive task and assesses whether it constitutes a malicious activity (Talebirad and Nadiri, 2023).

### v2

@@ 5.3.1. Defense on Knowledge Poisoning
(Baracaldo et al., 2017) proposes a new method for detecting and filtering poisonous data in the training sets of supervised learning models. It specifically utilizes data provenance to identify groups of data with a high correlation in their likelihood of being poisoned. This innovative approach aids in the effective identification and removal of malicious data.  (Yan et al., 2023) presents ParaFuzz, a novel framework for detecting poisoned samples at test time in LLMs, leveraging the interpretability of model predictions. The effectiveness of PARAFUZZ heavily depends on the specific prompts used with ChatGPT, which is employed to ensure high-quality paraphrasing. To optimize the detection process, the study adopts fuzzing to develop precise paraphrase prompts. These prompts are designed to effectively neutralize backdoor triggers while preserving the semantic integrity of the text.
There is still a significant gap in research focused on developing efficient defense strategies to protect LLMs from knowledge poisoning attacks (Das et al., 2024). Furthermore, empirical evidence indicates that LLMs are increasingly susceptible to these attacks. Current defense mechanisms, such as filtering data or reducing model capacity, provide only limited protection and often result in decreased test accuracy (Wan et al., 2023).
@@ 5.3.2. Defense on Output Manipulation
Moreover, inspired by human recursive thinking in the Avalon game,  (Wang et al., 2023b) introduces Recursive Contemplation (ReCon), a framework designed to enhance LLMs’ ability to detect and counter deceptive information. ReCon employs formulation, which generates initial thoughts and speech, and refinement, which improves these outputs. It also includes two perspective transitions, aiding LLMs in understanding others’ mental states and how others perceive their own mental states.
In addition,  (Xu et al., 2023a) has developed a benchmarking framework called MAgIC, designed to evaluate LLMs in multi-agent environments. It utilizes games and game theory scenarios to test models on reasoning, cooperation, and adaptability. The research employs Probabilistic Graphical Modeling (PGM) to enhance models’ capabilities in handling complex social interactions.
Finally, to address privacy concerns in LLM conversational agents,  (Bagdasarian et al., 2024) proposes AirGapAgent, which mitigates output manipulation by restricting access to only task-relevant data. It employs data minimization to prevent sensitive information exposure and introduces context isolation to prevent unauthorized access and manipulation. Additionally, a request escalation mechanism ensures user oversight in uncertain cases, balancing privacy and functionality.
5.3.3. Defense on Functional Manipulation
Given the emergence of functional manipulation as a new risk associated with the deployment of LLM agents, research on this specific threat remains limited. Thus, proactive security measures are essential. When using third-party LLM agents, it is crucial to protect personal privacy and be wary of excessive personal data requests by third parties. Users should limit data sharing, especially avoiding sensitive or personally identifiable information during interactions with LLM agents. Additionally, understanding and utilizing the data protection settings offered by LLM agents is vital. Adjusting privacy settings helps control what data can be collected and processed. Choosing providers with a strong reputation and transparency is also recommended, as these providers should have clear data usage and privacy protection policies along with a robust security track record (Zhang et al., 2024b).
To further address the challenges posed by functional manipulation, the introduction of the ToolEmu (Ruan et al., 2024) framework represents a significant advancement. This framework employs a language model to emulate tool execution, which allows for extensive and scalable testing of LLM agents across diverse scenarios and toolsets. Together with an LLM-based automatic safety evaluator, ToolEmu facilitates the identification and quantification of risks by examining potential failures and subsequent consequences. This method provides a dynamic alternative to traditional static sandbox evaluations, enhancing the ability to detect and mitigate high-stakes, long-tail risks effectively.
In addition, the Situational Awareness Uncertainty Propagation (SAUP) (Zhao et al., 2024) framework offers a novel defense mechanism by propagating uncertainty throughout the multi-step reasoning process of LLM agents. Unlike traditional methods that focus solely on single-step uncertainty, SAUP incorporates situational awareness by assigning weights to uncertainties at each step, enabling real-time detection of functional manipulation attempts. By identifying logical deviations and monitoring accumulated uncertainty, SAUP enhances the reliability of LLM agents and mitigates the risk of executing malicious operational patterns.
Finally,  (Anderljung et al., 2023) proposes an initial set of safety standards as an essential first step in industry self-regulation. These standards include pre-deployment risk assessments, external reviews of model behavior, the use of risk assessments to inform deployment decisions, and monitoring and responding to new information about model functionality post-deployment. This approach contributes valuable insights to the broader discussion on balancing public safety risks with the benefits of innovation in AI development.
@@ 5.3.3. Defense on Functional Manipulation
Table 3. Summary of Defensive Strategies Against Specific Threats
Threats Method Name Key Mechanism Performance Advantages / Limitations
Knowledge Poisoning Provenance-Based Poison Detection (Baracaldo et al., 2017) Utilizes data provenance to detect and filter poisonous data in training sets Superior in detecting data poisoning, enhancing model security and outperforming baseline defenses in adversarial settings Adapts to varied data trust levels, increasing flexibility / Requires substantial computational resources
ParaFuzz (Yan et al., 2023) Uses interpretability of model predictions to detect poisoned samples, employing fuzzing for precise paraphrase prompts Outperforms baseline methods like STRIP, RAP, and ONION across various datasets and attack types Excels against covert attack / High computational costs
Data Filtering & Reducing Effective Model Capacity (Wan et al., 2023) Utilizes data filtering to remove high-loss examples and reduces model capacity to hinder learning from poison data Lowers poison effectiveness from 92.8% to 21.4% and adversarial misclassifications to 35.2%, with a 3% accuracy drop Reduces poisoning effectiveness / Demands trade-offs between performance and safety
Output Manipulation BERTective (Fornaciari et al., 2021) Enhances BERT with additional attention layers to detect deception in Italian dialogues BERT enhances performance when combined with attention mechanisms to identify deception cues Enhances deception detection accuracy / Limited effectiveness of broader contexts
ReCon (Wang et al., 2023b) Employs formulation and refinement processes with perspective transitions to understand mental states ReCon boosts LLMs’ deception handling, increasing good side success from 15.0% to 19.4% in the Avalon game Enhances ability to discern and counteract deception / The dual-model architecture increases computational costs and complexity
MAgIC (Xu et al., 2023a) Uses games and game theory, combined with PGM, to evaluate LLM agents boosts LLM abilities by an average of 37% Enhances ability to navigate complex social and cognitive dimensions / Still in preliminary stages with limited scenarios
AirGapAgent (Bagdasarian et al., 2024) Employs context minimization, isolation, and request escalation to restrict data access Achieves 97% privacy protection with minimal utility loss, maintaining 88-90% task performance across various LLM models Provides strong privacy protection and maintains high utility / Depends on precise context definitions and may exclude essential data
Functional Manipulation ToolEmu (Ruan et al., 2024) Utilizes a LM to simulate tool execution and assess agent risks through an automatic evaluator ToolEmu offers precision rates of 72.5% and 68.8% with standard and adversarial emulators, respectively, while reducing setup time by 96.9% Offers flexibility and dynamic testing capabilities / Emulators may overlook essential constraints
SAUP (Zhao et al., 2024) Propagates uncertainty across all reasoning steps and integrates situational awareness for better reliability Improves AUROC by up to 20% across various datasets compared to baseline methods Offers comprehensive uncertainty estimation, strong compatibility / Depends on manual annotations, high cost, and limited generalization
Safety Standards (Anderljung et al., 2023) Proposes pre-deployment risk assessments, external reviews, informed deployment decisions, monitoring post-deployment Establishes regulatory frameworks, enhances risk assessments, and implements strict control measures Balances safety risks with innovation benefits / Needs further research and regulatory refinement
@@ 6. Future Trends and Discussion
With the advancement of LLM agents, their enhanced capabilities in complex observation, reasoning, and task execution have significantly broadened their application domains. Particularly, the development of Multimodal LLM (MLLM) agents enables processing of diverse data types, including text, images, and audio, further expanding their practical applications. Additionally, Large Language Model Multi-Agent (LLM-MA) systems support collaborative execution of sophisticated tasks. The integration of these technologies contributes to building more intelligent and efficient systems. Despite these advancements, significant privacy and security challenges have emerged. This discussion of future trends aims to provide insights for researchers, developers, and policymakers to optimize these technologies while addressing associated risks.
Figure 16 summarizes key development trends and associated security and privacy concerns of MLLM agents and LLM-MA systems, guiding the subsequent detailed discussions.
Figure 16. Overview of Development Trends and Security & Privacy Concerns
@@ 6.1.1. The Development of MLLM Agent
Recent advancements in LLMs have significantly surpassed traditional boundaries of language processing. These models now incorporate supplementary components such as instruction, interface, tools, knowledge, and memory, evolving into intelligent LLM agents that demonstrate expanded reasoning and expertise. Research studies  (Yang et al., 2023a; Wu et al., 2023) indicate efforts to bridge the gap between language models and multimodal tools, with intelligent agents like Visual ChatGPT  (Wu et al., 2023) and MMREACT  (Yang et al., 2023a) employing sophisticated prompt engineering techniques to achieve this target. Such efforts have given rise to the field of Multimodal Large Language Models (MLLMs). The general architecture of the MLLM is depicted in Figure 17.
@@
MLLMs extend LLMs with multimodal capabilities, enabling the processing of text, image, audio, and video. This enhancement facilitates a comprehensive understanding across these modalities (Xie et al., 2024). These models have found applications in medical imaging (Moor et al., 2023) and document processing (Liu et al., 2024c).
Furthermore, MLLMs have evolved into multimodal agents, such as embodied agents (Huang et al., 2024b) and graphical user interface agents (Wang et al., 2024b), which enhance their interactive capabilities in physical environments. These agents utilize MLLMs as planners, following natural language instructions and integrating perception, reasoning, planning, and execution capabilities to effectively operate in real-world settings (Xie et al., 2024).
MLLM agents are advancing towards AGI, enhancing their capability to understand and respond to complex human commands effectively.
@@ 6.1.2. The Security and Privacy Research on MLLM Agent
The development of embodied agents capable of interacting with the real world has become a highly active area of research. However, MLLM agents also present several security vulnerabilities, one of which is the phenomenon of multimodal hallucinations.
@@
Unlike language hallucinations, multimodal hallucinations refer to the phenomenon where the output descriptions generated by MLLMs are inconsistent with the actual content of images  (Yin et al., 2023), as shown in Figure 18. These phenomena manifest in two primary forms (Lee et al., 2024): one involves generated content that includes objects which are inconsistent with or absent from the target image  (Zhai et al., 2024; Liu et al., 2024a); the other, a more complex form, encompasses holistic misrepresentations of entire scenes or environments (Sun et al., 2023).
Similarly, adversarial attacks pose another critical threat to MLLM agents. These attacks involve crafted inputs that exploit the model’s vulnerabilities to produce biased or undesired outputs (Shayegani et al., 2023). For example, adversarial examples in multimodal settings may use subtle perturbations in images or audio to mislead the model into generating incorrect textual outputs or decision paths. Current methods to address these challenges include:
  • Hallucination Mitigation: Approaches such as utilizing self-feedback with visual cues to enhance model accuracy (Lee et al., 2024), employing instruction-tuning techniques to refine the model response to human instructions (Liu et al., 2024a), and implementing error-correction processes that identify and rectify hallucinations within the generated text (Yin et al., 2023).
  • Adversarial Defense: Techniques such as adversarial training (Yuan et al., 2024), data augmentation (Yin et al., 2024), and multimodal robustness frameworks (Gao et al., 2024) aim to improve the resilience of model against adversarial inputs.
Despite these advancements, significant gaps remain, particularly in systematically evaluating the effectiveness of these mitigation strategies and understanding the trade-offs between robustness, computational efficiency, and application performance. Future research could focus on the following:
  • Developing unified benchmarks and evaluation metrics to assess the robustness of MLLM agents against hallucinations and adversarial attacks.
  • Investigating the scalability of current mitigation strategies in real-world applications, such as healthcare and autonomous systems, where multimodal inputs are critical.
  • Exploring privacy-preserving methods, such as differential privacy, to ensure secure handling of multimodal data.
Improvements in the safety and reliability of MLLM agents require the development of robust mechanisms to detect and mitigate vulnerabilities. These advancements will ensure that MLLM agents can function securely and effectively in diverse real-world applications, paving the way for their broader adoption and ethical deployment in AI-driven technologies.
@@ 6.2.1. The Development of LLM-MA System
LLM agents exhibit advanced reasoning and planning capabilities, approaching human-like levels of decision-making and interaction. These agents are adept at perceiving their environments, making informed decisions, and executing actions based on complex contexts (Yao et al., 2023b).
@@
The application of LLM-MA systems spans across various fields, broadly categorized into two main types: problem solving and world simulation (Guo et al., 2024). For problem-solving applications, such as multi-robot systems (Mandi et al., 2023) and software development (Du et al., 2024), these systems enable interactions among diverse agents. This collaborative capability effectively solves complex real-world problems, mirroring the cooperative nature of human group work in tackling multifaceted challenges. On the other hand, world simulation encompasses applications such as society simulations (Park et al., [n. d.]) and game simulation (Wang et al., 2023b). These systems have demonstrated significant potential in various domains, showcasing their adaptability and efficiency.
@@ 6.2.2. The Security and Privacy Research on LLM-MA System
As research on LLM-MA systems increases rapidly, numerous challenges have emerged. Each agent within a multi-agent system may need to access and process sensitive data, and even execute code. Moreover, due to the intercommunication and interconnection between agents, security issues originating from a single agent can have profound and amplified effects in a multi-agent scenario. This has intensified the need for focused discussions on security and privacy issues in multi-agent environments.
One of the issues is hallucination, where agents generate outputs based on incorrect or fabricated information, representing a significant challenge for both LLMs and LLM agents. This problem becomes even more complex in a multi-agent context due to the interconnected nature of these agents and their frequent communication. Misinformation originating from a single agent can propagate across interconnected agents, creating a cascade of erroneous outputs throughout the system (Ju et al., 2024).
Another critical issue is the presence of malicious agents within the system. In one case, these agents may operate in a passive listening mode, where they receive information shared by other agents to perform tasks, but at the same time, they leak confidential information to attackers deliberately (Bagdasarian et al., 2024). In another case, malicious LLM agents may engage in an active communication mode, spreading virus-infected files, phishing messages, or other malicious code, attempting to attack or disrupt other agents within the system (Fang et al., 2024). Efforts to address these challenges have focused on two key areas:
  • Hallucination Mitigation: It is crucial to correct errors at the individual agent level and also to manage the flow of information between agents, thereby preventing the spread of inaccurate information throughout the entire system (Guo et al., 2024).
  • Malicious Agent Detection and Mitigation: Incorporating human feedback and user authorization for each step can help reduce these threats. This necessitates designing the system with robust security measures to prevent unauthorized access or misuse. An effective approach is the implementation of a stateless oracle agent, which can monitor each sensitive task and assess whether it constitutes malicious activity (Talebirad and Nadiri, 2023).
Despite their importance, the privacy and security challenges of LLM-MA systems remain underexplored in existing research. Future research could focus on:
  • Developing scalable and decentralized security solutions, such as blockchain-based communication protocols, to enhance inter-agent collaboration while minimizing systemic vulnerabilities.
  • Investigating privacy-preserving methodologies, including differential privacy and federated learning, to securely manage sensitive data shared among agents without compromising performance.
  • Creating standardized benchmarks and testing environments to systematically evaluate the robustness of LLM-MA systems against cascading misinformation and malicious activities.